07/02/21 Power appears to now be stable. All servers are running and plugged into there correct UPS boxes. I will leave Rex as intermittent overnight incase we have any short outages from UK Power Networks. My thanks to all customers for your patience today. 17:16
07/02/21 Our external power is now completly severed and completion has been moved back to between 16:00 and 17:00 this afternoon. This has not impacted the servers in any way. 14:22
07/02/21 Rex server now up and running. Power issues should be resolved late this afternoon. At that point the servers will be shut down and re-started in an orderly manner. I don't know at what time. However the outage in services should not be longer than approx 15 minutes if that. 12:16
07/02/21 Sub station running low voltage this am Backup generator had a fuel flow problem now resolved. Rex web and e-mail still offline. All other services now recovering. 10:57
07/12/20 All ADSL and FTTC services have been returned to normal over the weekend. 08:55
04/12/20 Our wholesaler has a UK wide rolling outage which is affecting some people but not all. They are moving affected customers around there network so that they have access. This involves losing internet for 10 mins or so. After that a re-boot of the router may be required. Once the problem is cleared I will post that here. 16:00
23/11/20 Most internet services have been working this afternoon. Our wholesaler will continue monitoring for the next 24 Hrs.
23/11/20 A temporary fix has been implemented to restore service while BT continue to investigate the cause of the outage. 13:50
23/11/20 Our internet wholesaler has a major aoutage which started at 12:29 today which is a UK wide issue. It is being investigated. This is affecting a good proportion of our internet customers and we lost our own connection for about 5 mins. Service will be spotty this afternoon but get better as time passes. 13:15
09/11/20 It is with regret we annouce the end of the zfc.org.uk free e-mail accounts on the 12th November 2020. I do hope that all the users over the years have enjoyed the friendships from this.
27/10/20 Our earlier incident actually turned out to be a customers e-mail account spamming. This has been resolved and I would mentioin that the password used was NOT one we had set it was far to weak hence the incident. PLease if you change the password to any account with Orpheus make it a strong one. This will help us runa safe and good service for you all.
22/10/20 We made a small change to the way mail was routed for one of the servers on Monday. It has since become apparent that this has caused more problems than it solved. The servers have been set back to the way they were. Various customers have had mail deferred by Verizon because of the change this should correct that situation but it may take 24 hours or so and I don't know what may happen to the mail that has stalled. Please if possible re-send these mails either on a diffent account or wait a couple of days and re-send them.
04/09/20 Many customers have asked after my family in recent months and know that my daughter Serena has been furloughed and working with me on our accounts. I would like to thank her and also let you all know that she has been able to retain her job at Westminster Abbey.
19/08/20 We will have 2 service disruptions in the coming week. All servers will be shutdown at some point between 10~12 am on Saturday to allow connection of a backup generator and testing. On Wednesday 26th August a powerline pole re[placement will occur between 10~12am I hope that there will be minimal service loss on Wednesday or possibly none at all.
03/08/20 After a really hectic time since the start of the lockdown I am taking some time out to do some renovations at home. So from Wednesday 5th Aug to Tuesday 11th Aug I will be away from the office. The phone and e-mails will still work so please leave a message but understand it may take me an extended time to reply. 15:55
24/07/20 If anyone still has an issue from yesterday please power down your router and allow 10 minutes then power back up. Any further issues please get in contact in the usual ways. 19:24
23/07/20 I have just been informed that there has been a fire in a BT major Node building which has caused this issue. BT are working hard to re-route services and we are seeing a gradual return to the disrupted services. I cannot say when any customers specific line will return but it will be asap as BT continue there work. We do hope that no one was injured during the fire. 14:24
23/07/20 We are aware of a MSO in the Northern UK which is affecting customers internet access. We currently do not know what or when it will be resolved. We will post this info up once we know. 12:56
17/07/20 Pleased to say no issues with Rex. But I now have a better understanding of some of our security. 09:10
16/07/20 We may have an issue with our Rex server so a re-start may need to be done during work hours this normally takes about 10 mins before all services re-start. If you notice an issue today this will most probably be why. 12:25
11/06/20 Our server man Dan has done a great job and most services are now running smoothly. Thankfully this level of severity doesn't occur to often.
10/06/20 It appears we have been having some 'extra' Brute force attacks in the last 24 hours which has impacted some but not all services. These have had patchy effects to our services in some instances and longer duration ones on others. We are looking into it to see how we can mitigate and stop the trouble. I am sorry for any issues this has caused our customers.
04/05/20 Last reset on the server was late yesterday afternoon. We now expect that it should run properly. We will monitor this as we start to setup the new version and run that as a test server before we transfer all the current accounts across. We will inform everyone before we do this so you are all aware.
03/05/20 We do have a service running at the moment on the smtp2 server. Work hasn't been finished yet so service may be erratic over the next day or so. I will update here with news as soon as I can. 00:52
02/05/20 Work is currently going on with the smtp2 server. This is interupting service so I would suggest that for the next couple of hours if possible no-one sends mail using this server. I will post an update later on when we stop and the server is stable. I am sorry for the inconvenience this is causing you. 19:21
02/05/20 Smtp2 is running. We will be fine tuning the server over the next couple of days. I am leaving the status light as intermittent because of this. We will do our best not to interrupt service but we will need to do the occasional restart as we work. Hopefully this will not cause you any problems. 12:00
02/05/20 Smtp2 teething problems with the new firewall software this am. 09:30
01/05/20 The replacement server is in place and as far as I have tested this morning appears to be working. We transferred over all the accounts from the old server so if you have an issue please contact us and we can refresh your password etc.. to get you back to normal. My grateful thanks to Dan for working so long and hard for us yesterday on this issue. 09:15
30/04/20 Last update for tonight as of 19:30 we started to re-build the smtp2 server from scratch not our new version but a fresh version of the current server. By doing this we hope to avoid the repeated failure we have seen today. I anticipate that there will be no service tonight and possibly well into Friday. I do understand in these times how frustrating this is. Once I have been updated I will see if I can sort out a different route for you, if it is urgent. But it will entail changing your mail settings and then changing them back once the smtp2 is restored. I will update you in the morning on that. 20:20
30/04/20 smtp2 is back down again and we don't know why this is failing. I would suggest if you have your broadband with Orpheus that you change the send server to use smtp.orpheusnet.co.uk and set up according to your original account details. I am currently still in the office if you have questions about the setup. We will do out best to sort this out but I think this may take us a while to sort. 16:51
30/04/20 Smtp2/sarge server is now back up and running with all users restored as of 14:30 Also plans are being made to replace the server in short order. It was on the road map but is now at the top of the list. 15:05
30/04/20 Restoring the server is progressing but has taken much longer than it should due to some tech issues which are now resolved. I will post up when all is working correctly. 12:00
30/04/20 Issue with smtp2 this am we are currently in the process of restoring the backup. 09:37
28/04/20 We are receiving some reports of slow response times on various servers which are not on the same physical hosts. We cannot currently find any reason for this all systems are running nominally. We are investigating to see if there is an outside issue, but again currently can see nothing wrong.
22/04/20 FTP on Rex is working. An errant website was consuming all the CPU this site is now suspended and the customer will affect repairs with our assistance. 13:09
22/04/20 FTP access to websites on Rex currently is not working. Next update in approx 90 mins. 10:45
18/04/20 We have persuaded the server to behave so all services on Rex should now be getting back to normal. Sarge/smtp2 is also now running and up to date. My thanks for your patience. 19:57
18/04/20 We have had 'an advised' uexpected power outage this evening. On restart the server we had issues with last night is being difficult this runs Rex which does website and e-mail for some customers so you are currently affected. Sarge as ever causes us problems but that is being downloaded from our backup and should be up within the next 45 mins or so. This server hosts the smtp2 authorised send server. 19:10
18/04/20 New link to this news page has now been started 15:12
18/04/20 All services have now been fully restored. 11:15
18/04/20 vmHost server went down during the night currently restoring the vm's. 06:30
27/02/20 rDNS/PTR issue at our suppliers servers has now been resolved and is propagating out as I write this. Services should recover fully within the next 24 hours. 15:47
27/02/20 Over the past 24Hrs or so various services including e-mail have had odd issues. It appears that the reverse DNS that is set at our suppliers is not reporting correctly for odd IP numbers but is for even ones. This was reported during the morning and they are investigating urgently into the matter. This is unfortunately totally outside of our control and all we can do is wait for the investigations and resolution to happen.
17/02/20 All services have now been fully restored as of 08:00 this am. My thanks to the technicians up and down the country who worked hard through the night for all of us affected by Storm Dennis.
17/02/20 Due to storm damage locally we lost power and access from approx 5:30pm to 11:45pm on Sunday. All services except smtp2 are back and working. Smtp2 couldn't be fixed as our server tech had no access to the servers he will login on Monday am to look at this service. I don't expect that it will take too long to sort.
16/02/20 smtp2 service server has issues today since mid-morning. We are working on returning the service back to normal. The weather affecting the UK will not be helping us all.
02/02/20 We have reports of a Phishing email this morning badly written incorrect email address all the usual signs that it has nothing to do with Orpheus. Please do not click on any links just delete the mail. I will be adding a page specifically for this pernicious type of scam on our support pages in the next few days with hints and tips, how to best keep yourself safe.
15/01/20 Issue resolved at 07:54 this morning.
14/01/20 smtp2/Sarge is currently unavailable. The service should be restored during the morning once the server tech is onsite. My apologises for the inconvenience this will be causing.
06/01/20 Wednesday night we upgraded the Roundcube Webmail system on the 'new' web server Rex. Please get in contact if you would like me to upgrade you. You can of course do the upgrade yourself by logging into the control panel as your administrator going to services/web applications and clicking on the little pencil icon on the line for the webmail client.
03/01/20 We have reverted to a backup copy of Sarge/smtp2 and the service has been restored. Going forward we will need to replace this server with a new authenticated server we will now begin this program. 11:51
03/01/20 Sarge/smtp2 was updated and re-started early this morning. At this time sending using smtp2 has not been possible as at 09:55 our server tech will actively be looking at the server. Updates to follow. 09:50
Net Station compatibility - We've now acquired the capability to re-program NC/set-top box smart cards in order to connect to Orpheus. If you would like us to re-program your smart card, please return it to us together with a stamped addressed envelope and a cheque for six or twelve months subscription. We will setup a new account and return your smart card, reprogrammed to connect to Orpheus Internet. (This service is no longer available 2009)